OpenStudy (anonymous):

Solve the following system of equations. -2x + 5y = 60 x + 5y = 45

OpenStudy (uri):

-2x + 5y = 60 x + 5y = 45 - - - =========== -3x=15 x=-5

OpenStudy (uri):

x + 5y = 45 -5+5y=45 5y=45+5 5y=50 y=10

OpenStudy (lieutenantgeneral):

-2x + 5y = 60 x=5n y=2(n + 6) x + 5y =45 x=5n y=9 - n

OpenStudy (anonymous):

thank you @LieutenantGeneral @uri

OpenStudy (lieutenantgeneral):

You Welcome